OpenStudy (anonymous):

HELP!! What choice is equivalent when x>0? √98x^3 divided by √72x^2 1) 7x/6 2) 7√x/6 3) √98x*3-72x^2 4) √26x

OpenStudy (anonymous):

dude rewrite that and use parenthesis

OpenStudy (anonymous):

sqrt(98) = sqrt (49 x 2) so you can pull out the 49 as 7, sqrt(72) = sqrt (36 x 2) so you can pull out the 36 as 6, similarly, the x^3 can be pulled out as x^2 and the x^2 can be pulled out as x. this would leave you with (7x^2 times sqrt(2)) over (6x times qrt(2)) cancel out and move from there.

OpenStudy (anonymous):

7x/6 is my answer

OpenStudy (anonymous):

7x/6 is my answer

OpenStudy (anonymous):

*nods*

OpenStudy (anonymous):

thanks burton! that really helped me on how to solve this . :)

OpenStudy (anonymous):

do I not get a point for checking his work?

OpenStudy (anonymous):

thanks guys***

OpenStudy (anonymous):

because if x = 1 then then its root(98)/root(72) = root(2*7*7)/root(2*6*6) = 7root2/6root2 =7/6 with x=1